OpenStudy (kwame):

In a certain school, 40% of all students are 8 years old or younger. The number of the remaining students is 120. How many students are 8 years old or younger?

OpenStudy (rsmith6559):

What percentage of the students are > 8?

OpenStudy (bluprints):

It's just like, what is the 40% of 120. Equation is Y=P% * X where: Y=value P%=40% X=120 *(percentage of all students who are 8 years old or younger)40%--> 40/100(to convert percentage to decimal)= 0.4 *120 (remaining students) *120*0.4=48 So there are 48 students who are 8 years old or younger. I hope it helps.

OpenStudy (sudhansharma):

let the no. of students are ; x- 8 years old y- more than 8 z- less than 8 then 60/100*(x+y+z)=120 so, x+y+z=200 now, (x+z)/(x+y+z)*100=40 so x+z=80 so, 80 students are 8 years or younger.